Please answer ALL of the parts! Thank you and hope you're staying safe :)

A)

Consider the following reaction:

PCl5(g) PCl3(g) + Cl2(g)

If 7.65×10-3 moles of PCl5, 0.257 moles of PCl3, and 0.284 moles of Cl2 are at equilibrium in a 18.1 L container at 616 K, the value of the equilibrium constant, Kp,

is __________.

B)

Consider the following reaction:

2HI(g) H2(g) + I2(g)

If 1.44 moles of HI, 0.258 moles of H2, and 0.211 moles of I2 are at equilibrium in a 15.2 L container at 883 K, the value of the equilibrium constant, Kp, is __________.

C)

The equilibrium constant, Kp, for the following reaction is 1.80×10-2 at 698K.

2HI(g) H2(g) + I2(g)

If an equilibrium mixture of the three gases in a 19.4 L container at 698K contains HI at a pressure of 0.348 atm and H2 at a pressure of 0.756 atm, the equilibrium partial pressure of I2 is __________ atm.
Answer #1

cone = mols r. Pus] =7.65*10 mol = 4.23 810-4M 18.1L [olg] = 0.284 mol = 0.0157 M 18-1L Tile 2] = 0.257 mol = 0.0142 M 18.12